Calculate the mass of Ethylene glycol which is required to prepare thousand Graham aqueous solution with a difference of 10 degree Celsius and its freezing point km equals to 1.86 degree Celsius kg per mole

Grade:12

1 Answers

Vikas TU
5 years ago
Freezing point depression formula is given as:
Tf = iKf*m
10  = 1*1.86*m
m = 5.37
moles = m*1000 => 5370
weight = 5370*molar mass
          = 5370*62
          = 332940 g